I cannot figure this out. Can someone please show me how?
3/x - 1/6 = 1/x
Three over x minus 1 over six = 1 over x
Thank you!

Answer by checkley71(8403)

3/x-1/6=1/x or 3/x-1/x=1/6 or 2/x=1/6 or x=12
proof 3/12-1/6=1/12 or 3/12-2/12=1/12 or 1/12=1/12